[figure: 9.7 cm]

Electronic Structure and Magnetic Properties of (Y_{1−x}Gd_{x})_{n+1}Co_{3n+5}B_{2n}

P. Vlaic^{a}, E. Burzo^{b}

a Biophysics Department, University of Medicine and Pharmacy “ Iuliu Hatieganu” ,

3400 Cluj-Napoca, Romania b Faculty of Physics, Babeş-Bolyai University, RO-3400 Cluj-Napoca, Romania

Magnetic measurements, in the temperature range 4.2-1000 K, and band structure calculations were performed on (Y_{1−x}Gd_{x})_{n+1}Co_{3n+5}B_{2n} compounds with n=0; 1; 2; 3 and ∞. The experimental magnetic moments per formula unit are well described by the computed values. The substitution of Y by Gd increases gradually the magnetic moments on cobalt sites. In addition, the polarizations of Y (4d) and Gd (5d) bands increase. The 4f-5d-3d and 4d-3d exchange interactions are analyzed. Starting from the position of the center of gravity of Co (3d) bands, a model is proposed in order to take into account their magnetic moments in correlation with their local environments.

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : poster

Corresponding author : P. Vlaic

Address for correspondence : Biophysics Department University of Medicine and Pharmacy “ Iuliu Hatieganu” 3400 Cluj-Napoca ROMANIA

Email address : petrica vlaic@yahoo.com